Science Report – March 23rd

Crew 207

Research Report

Submitted by Crew Scientist Tiffany Ni

3D Printing Projects

Today is the last full day of MDRS for myself and our crew HSO. Our 3D-printing experiments continue to progress. We have completed printing all of our labware, International Space Station medical inventory items, STEM teaching aids, and are now printing some things here and there around MDRS, as well as other miscellaneous 3D printed projects.

To date, we have printed:

1. Western blot combs (18 well combs for making gels)
2. Microscope slide holder (reusable holder for square coverslips)
3. Sieve mesh (2.0mm holes)
4. Sieve structure (for holding the mesh net)
5. Mobile phone microscope clip
6. Sample box (for 96 tubes)
7. Stackable freezer bin
8. Salt spray chamber drain
9. Triangular straws (in a rainbow of colors)
10. Sieve mesh (5.0mm holes)
11. Sieve structure (for holding the mesh net)
12. Slide box holder (for storing microscope glass slides)
13. Test tube rack
14. Test tube holder
15. Sample rotator components
16. Small test tube holder (for the freezer)
17. Heat gun nozzle mount
18. Tongue depressors (in a rainbow of colors)
19. Stoma bag clip (used as a chip bag clip)
20. Full set of fraction tools (mathematics educational tools)
21. Funnels (small, medium, large)
22. Construction blocks
23. 12 tube separator (for Eppendorf tubes)
24. The side part of thoracentesis trainer
25. Customized finger splint (originally printed on the ISS0

During our time at MDRS, our 3D printed straws proved incredibly useful for our crew members. It has helped greatly in our attempts to reduce our water consumption. At MDRS, we have also demonstrated that a number of our 3D printed labware items are functional – as validated by two basic science research students.

Even though we will be leaving tomorrow morning, we will continue to stay in contact with our MDRS crew members and send 3D printable files for incoming crew members to print! (Perhaps some reusable bubble tea straws?)

Signing off,
Tiffany Ni
Crew 207 Scientist

EVA Report – March 23rd

Crew 207 EVA Report 23-Mar-2019

EVA #6

Author: Diane Rothberg

Purpose of EVA: Drone map areas that will be used to create a 3D printed map of the MDRS Hab area.

Start time: 10:00
End time: 12:00

Narrative: Two Part 107 drone pilots mapped two areas outlined in Drone Deploy software. First one near Hab. Second one climbed the hill near Pooh’s corners and launched the drone from there.

Destination: Hab area, and Pooh’s Corners

Coordinates 4255000 x 518000, and 4251000 x 518920

Participants: Dean Jin, Tom Baldwin, Diane Rothberg

Road(s) Route 1103 near the Hab and Pooh’s Corners area

Mode of travel: Walking and Rovers (Opportunity + Curiosity)

Crew 207 EVA Report 20 Mar 2019

EVA #7

Author: Julielynn Wong

Purpose of EVA: Allow novice drone pilots to compare a standard versus novel controller to fly a camera drone in a square test flight pattern.

Start time: 14:40
End time: 16:40

Narrative: The weather was clear and sunny.

Two EVA Team members flew the traditional controller first for a test flight in a square configuration (60 ft for each side) at an altitude of 20 feet. Then the EVA team member flew a novel controller for the same square flight test pattern. XO also use the novel controller to conduct a habitat roof inspection. A post-flight survey was completed by all EVA team members.

Destination: MDRS sign

Coordinates: 4250850 / 518100

EVA Participants: Dean (XO), Diane Rothberg (GHO), Tiffany Ni (SCI)

Roads and routes per MDRS Map: Road 1103

Mode of travel: Walking

Greenhab Report – March 23rd

Crew 207 GreenHab Report – 23 Mar 2019

GreenHab Officer: Diane Rothberg

Environmental control: Heat and cooling w/ambient air 4:50 hr

Shade Cloth (40%)

Average temperatures: RH

Low: 16.0 deg C 20%

High: 28 deg C 42%

Hours of supplemental light: 0 Hours

Daily water usage for crops: 6 gal

Water in Blue Tank: 148 gallons

Crops watered at 07:30 hours (3 gal) and 16:45 hours (3 gal)

TIME NOTED TEMPERATURE (DEG C) / HUMIDITY (%)

1) 7:30 AM 18 deg C / 37%
2) 12:00 AM 30 deg C / 20%
3) 14:45 PM 24 deg C / 20%
4) 16:45 PM 23 deg C / 23%

Changes to crops: Radishes planted 3/20 showing small sprouts

Narrative: None

Harvest: Basil: 6 gm
parsley: 4 gm

Support/ Supplies needed: none

Operations Report – March 23rd

Crew 207 Operations Report 23 Mar 2019

SOL: 7

Name of person filing report: Tom Baldwin

Non-nominal systems: Generator: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator:

Generator (hours run): (TBD, attempted contact with Dave.) hours

Generator was turned off at (TBD, attempted contact with Dave), SOC (TBD) %

Diesel Reading – 50 %

Propane Reading – 62 %

Ethanol Free Gasoline – 2 gallons.

Water (auxiliary tank) – Not in use

Water (static tank) – 160 gallons

Auxiliary to Static tank transfer– Not applicable

Gallons transferred: Not Applicable

Water in Green Hab – 148 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 142050.1 gallons

Toilet tank emptied: No

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Not used

Hours: 85.1 hours

Beginning charge: 100 %

Ending charge: 100 %

Currently charging: Yes

Opportunity rover used: Not used

Hours:54.9hours

Beginning charge: 100 %

Ending charge: 95 %

Currently charging: Yes (now 100%)

Curiosity rover used: Used

Hours: 84.7 hours

Beginning charge: 100 %

Ending charge: 97 %

Currently charging: Yes (now 100%)

Notes on rovers: not used

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): N/A

Reason for use: Not used

Oil Added?: N/A

ATV Fuel Used: N7A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar: None.

CrewCar used and why, where? David reports that Crew Car is back in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EVA suits normal, with all batteries greater than 13 volts, (including the three used for EVAs today).

Summary of Hab operations: Checked SOC of the electrical system.

08:00 79%, generator on, fault light on. Reported to David.
12:04 92%, generator off, fault light off.
16:00 74% generator off, fault light off. Reported to David.

Summary of Green Hab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Questions, concerns and requests to Mission Support: None

Sol Summary – March 23rd

Summary Title: Day #7 of Sim

Author: Julielynn Wong

Mission Status: All nominal

Sol Activity Summary:

Yesterday evening, COM continued data collection for the thoracentesis study.

This morning’s EVA successfully completed two additional drone mapping flights of MDRS. This afternoon’s EVA completed flight testing of a square (60 feet) flight pattern at an altitude of 20 feet using a traditional and new drone controller by two drone pilots. The third pilot conducted a habitat roof inspection with the new controller.

A post-flight survey was completed.

HSO + SCI continued 3D printing items in Level 1 of the Hab and tested their 3D printed labware in the Science Dome.

Look ahead plan: Clean up MDRS this evening for inspection. We will break sim tomorrow and transport two crew members (HSO + SCI) to GJT on Sol 8. ENG + GHO to depart on Sol 10. Mission summary to follow later on this week.

Anomalies in work: None

Weather: Sunny

Crew physical status: Fine

EVA: AM: Drone mapping PM: Drone controller testing

Reports to be filed: Operations, GreenHab, Journalist, EVA Report, Research Report

Support Requested: None

Journalist Report – March 23rd

Journalist Report March 23, 2019

Journalist: Diane Rothberg

A busy last day of simulation with two EVAs, both of which were similar to yesterday’s.

This morning, we had a great time stumbling up a hill near Pooh’s Corners with our expensive drone and controller, trying not to destroy them. But we were rewarded with a perfect launch site and a gorgeous view. Our camera drone flew its programmed route well and allowed us to keep it in sight at all times.

Tiffany and Kevin were busy in the Science Dome testing out the 3D printed labware they printed at MDRS using biodegradable plastic. The good news is that our colorful devices worked and looked fabulous!

During this afternoon’s EVA, Tiffany got a chance to fly a drone for the first time, and we both performed the control stick comparison. No contest! The new controller is much more intuitive than the traditional controller, and made a clumsy newbie like me feel like a “drone virtuoso!”

After dinner, we are wrapping up the thoracentesis data collection for our crew, and cleaning up MDRS for inspection.

Our last dinner together will be quite ethnic… potato latkes and lentil soup, concluding with one of Tiffany’s excellent desserts…. apple pie “a la Mars.”